OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3)with a slope of -6. Answer y = -6x + 15 y = -6x - 20 y = -6x + 20 y = -6x - 15

jimthompson5910 (jim_thompson5910):

y = mx+b -3 = (-6)(-2)+b -3 = 12+b -3-12 = b -15 = b b = -15 So the equation of the line is y = -6x-15